Create a family budget plan. Make sure to cover that 360 degrees in a circle and is used when calculating the angle measurement. Explain how you applied your knowledge about central angles and areas of the circle in creating a graph.


1
Expert's answer
2021-01-14T15:47:42-0500


The central angle is a part of "360\degree"

Let "\alpha" be central angle, "S_\alpha" be a square inside central angle, "L_\alpha" be a length of arc of central angle,

"S" be a square of circle, "L" be a length of circumference. Then:

"\frac{\alpha}{360\degree}=\frac{S_\alpha}{S}=\frac{L_\alpha}{L}"


In our case every, part of budget is represented by central angle with appropriate proportion.